# Current validation practice undermines surgical AI development

### v1

@@ 4.3. Experimental Design
Weighted phase-wise aggregation: Performance was first aggregated for each phase, and the resulting phase-level scores were then combined into a global score using clinician-defined weights reflecting the clinical relevance and complexity of each phase. For simplicity, we adopted a straightforward weight assignment strategy, in which phases deemed less relevant or complex were assigned a weight of 1 (phases 0, 4, 5, 6, 9, 12), intermediate phases were assigned 2 (phases 2 and 10), and highly relevant or complex phases received a weight of 3 (phases 1 and 8) [maier2021heidelberg].
Across all aggregation strategies, the aggregation operator can vary, e.g., it could be the mean, median, or other percentiles. In line with the original challenge, we used the 5th percentile to reflect worst-case performance. To assess the agreement between the default (frame-wise) ranking scheme and alternative aggregation strategies, we calculated Kendall’s tau [kendall1938new], a rank correlation coefficient, quantifying the similarity between two orderings (Kendall’s tau = 1: perfect agreement, Kendall’s tau = 0: no association, Kendall’s tau = -1: complete disagreement).
@@ 6. Acknowledgements
A.R.: This work was initiated by the Helmholtz Association of German Research Centers in the scope of the Helmholtz Imaging Incubator (HI). L.M.-H.: This project was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (NEURAL SPICING, 101002198), the National Center for Tumor Diseases (NCT), Heidelberg’s Surgical Oncology Program, the German Cancer Research Center (DKFZ). E.C.: This publication was further supported through state funds approved by the State Parliament of Baden-Württemberg for the Innovation Campus Health + Life Science Alliance Heidelberg Mannheim. O.C.: The research leading to these results has received funding from the French government under management of Agence Nationale de la Recherche as part of the “France 2030” program (reference ANR-23-IACL-0008, project PRAIRIE-PSAI), as part of the ”Investissements d’avenir” program (reference ANR-19-P3IA-0001, project PRAIRIE 3IA Institute and reference ANR-10-IAIHU-06, project Agence Nationale de la Recherche-10-IA Institut Hospitalo-Universitaire-6) and from the European Union’s Horizon Europe Framework Programme (grant number 101136607, project CLARA). R.D.: This project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101092646. D.D. is funded by NIH K23 EB034110. Q.D.: This work described in this paper was partially supported by a grant from the ANR/RGC Joint Research Scheme sponsored by the Research Grants Council of the Hong Kong Special Administrative Region, China and the French National Research Agency (Project No. A-CUHK402/23). G.F. is supported by Canada Research Chair in Computer-Assisted Surgery. S.Gia. is supported by the Royal Society URF R 201014. T.H. is a Consolidator Researcher, receiving financial support from the Distinguished Researcher program of Óbuda University. His work has been partially supported by ACMIT (Austrian Center for Medical Innovation and Technology), which is funded within the scope of the COMET (Competence Centers for Excellent Technologies) program of the Austrian Government. D.A.H. is supported by a grant from the American Surgical Association Foundation. F.R.K. receives support from the German Cancer Research Center (CoBot 2.0), the Joachim Herz Foundation (Add-On Fellowship for Interdisciplinary Life Science), the Central Indiana Corporate Partnership AnalytiXIN Initiative, the Evan and Sue Ann Werling Pancreatic Cancer Research Fund, and the Indiana Clinical and Translational Sciences Institute (EPAR4157) funded, in part, by Grant Number UM1TR004402 from the National Institutes of Health, National Center for Advancing Translational Sciences, Clinical and Translational Sciences Award. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. J.L.L. received funding by the Swiss National Science Foundation (P5R5PM 21766), Novartis Foundation for medical-biological Research (#23C162), and the Vontobel Foundation (0867/2024). H.J.M. is supported by the NIHR UCLH/UCL Biomedical Research Centre. N.P.: This work has received funding from the European Union (ERC, CompSURG, 101088553, PI: Nicolas Padoy). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work has also been supported by French state funds managed within the Plan Investissements d’Avenir by the ANR under reference ANR- 10-IAHU-02 (IHU Strasbourg). H.R.: Hong Kong Research Grants Council Collaborative Research Fund under Grant CRF-C4026-21G. S.S. is supported by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence ”Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Dresden University of Technology. D.S. is supported by the Royal Academy of Engineering Chair in Emerging Technologies. M.W.: This work has been funded by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy—EXC 2050/1—Project ID 390696704—Cluster of Excellence ”Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of TUD Dresden University of Technology and by the German Federal Ministry of Research, Technology, and Space within the “Surgical AI Hub Germany” project (grant number BMBF 02K223A110).
@@ 7. Conflicts of Interest
M.S.A. is a speaker of Medtronic and advisor of Johnson & Johnson. D.A. is a shareholder and has a leadership role in Scialytics. N.W.C.: Nil. O.C. reports having received consulting fees from Therapanacea (2022-2024). OC reports that other principal investigators affiliated to the team which he co-leads have received grants (paid to the institution) from Sanofi (2020-2022) and Biogen (2022-2023). OC reports that his spouse was an employee of myBrainTechnologies (2015-2023) and is an employee of DiamPark. G.F. declares the following conflict of interest: Johns Hopkins University, Harvard Brigham and Women’s Hospital. S.G. declares the following competing financial interests: he has or has had consulting relationships with Una Health GmbH, Lindus Health Ltd., Flo Ltd, ICURA ApS, Rock Health Inc., Thymia Ltd., FORUM Institut für Management GmbH, High-Tech Gründerfonds Management GmbH, DG SANTE, Prova Health Ltd, haleon plc and Ada Health GmbH and holds share options in Ada Health GmbH. T.P.G. is the Founder of SST. D.A.H. is a consultant for Medtronic. A.J. is an employee of Intuitive. F.R.K. declares advisory roles for Radical Healthcare, USA; and the Surgical Data Science Collective, USA, and has received research funding from Novartis. H.J.M. is employed by and holds shares in Panda Surgical Limited. P.M. is a co-founder and shareholder of Scialytics. H.N. is affiliated with Proximie Ltd, London, UK. N.P. is co-founder and own shares in Scialytics SAS. N.R. is an employee at NVIDIA. M.M.R. is now working at SAP Fioneer. D.S.: Medtronic, Odin Vision, Panda Surgical, EnAcuity, Helico Medical, Uncovr, Vope Medical. M.W. has received honoraria by KARL STORZ SE & Co KG.
@@
| SUPPL. NOTE 3 | Results of the systematic review per pitfall | SUPPL. NOTE 3 |
@@
| P1.1 | Non-representativeness, low quality, and low relevance of data | The data used for training and/or testing does not accurately reflect the intended real-world use case or lacks clinical relevance. This may, for example, result from geographical imbalance, a lack of diversity in patient demographics, regional differences in surgical standards, or limited or outdated variations in surgical techniques. Additionally, data selection criteria may exclude challenging but clinically relevant cases, such as poor camera quality, low lighting conditions, challenging procedures, or imaging artifacts, which are crucial for assessing model robustness in realistic scenarios. The inclusion of irrelevant data, such as out-of-body frames or unintentionally recorded segments, may also dilute the training signal or introduce misleading patterns. The use of simulated or experimental data may further reduce real-world applicability, while proxy data can misalign validation with clinical objectives. Variability in video quality and preprocessing protocols across institutions may further compromise data fidelity and model reproducibility. | [godau2023deployment, kolbinger2024strategies, koch2024distribution, sellner2023semantic, kilim2022physical, castro2020causality, subasri2025detecting, lavanchy2024challenges, kassem2022federated, mascagni2022multicentric, kong2021accurate, junger2022state, bar2020impact, paranjape2023cross, engelhardt2018improving, sahu2021simulation, huaulme2023peg, kolbinger2023anatomy, kolbinger2024artificial, ding2022carts, childers2021same, jogan2024quality] | Extended Data Fig. 1 |
@@
| P1.2 | Limited sample size/test cohorts, or data imbalance | The test data may be limited due to small sample size and/or the absence of validation on independent datasets, including those from different institutions, time periods, or populations (external validation), or from real-world clinical use after deployment (post-deployment validation). Limited data can mislead model development (e.g., algorithm selection), increase the risk of overfitting, or yield unreliable model performance validation, making it difficult to draw meaningful conclusions about the model’s real-world performance. These issues are amplified in the presence of class imbalance, where rare classes may appear in only one subset or fold, or be entirely absent, leading to inconsistent model performance depending on the specific validation or test set used. | [bar2020impact, kostiuchik2024surgical, yamlahi2023self, demir2023deep, killeen2023pelphix, maqbool2020m2caiseg, cheng2023deep] | Extended Data Fig. 2 |
@@
| P1.3 | Existence of spurious correlations within data | Spurious correlations are statistical patterns in the data that do not reflect robust or generalizable associations with the target task. They can arise from biases in data collection, labeling, or contextual factors unrelated to the underlying clinical objective, and often reflect dataset-specific artifacts that fail to generalize. This may lead to short-cut learning, where a model exploits irrelevant patterns, for example, achieving seemingly good performance by associating certain surgical outcomes with specific recording conditions, such as particular camera types used at different hospitals, or inferring the presence of instruments from irrelevant features such as glove color, if such features are biased within the dataset. | [yuan2024advancing, banerjee2024bias, mahmood2021detecting, geirhos2020shortcut, dammu2023detecting, bykov2023finding] | Extended Data Fig. 3a |
@@
| P1.4 | Incomplete annotation or missing contextual information | Annotations of the data are – intentionally or unintentionally – incomplete, leading to gaps in the information available for reliable validation. This may include missing metadata, partially annotated video sequences, the omission of specific objects, events, or surgical phases, or insufficient detail for the intended task. Missing metadata, such as age, for example, may lead to undetected confounders and thus biased models. Partially annotated video data may render the application of temporal metrics (e.g., for assessing continuity or stability across frames) infeasible and thus limit the ability to assess performance over time, such as continuity or stability across video frames. While some approaches may use partial annotations intentionally (e.g., in weak supervision), unmanaged or undocumented incompleteness may still compromise reliability and validation validity. The omission of relevant entities can result in inaccurate performance estimates, for example, underestimating sensitivity when relevant instances are not labeled, and may give a misleading impression of model accuracy and clinical utility. | [pangal2021guide, ward2021challenges, tajbakhsh2020embracing, twinanda2016endonet, funke2019video, peng2024reducing, nyangoh2023systematic, murali2023endoscapes] | Extended Data Fig. 3b |
@@
| P1.5 | Unreliable or inconsistent annotation | Unreliable or inconsistent annotations can undermine the validity of model validation by introducing ambiguity, bias, or uncontrolled variability. This may result from the lack or insufficiency of a standardized annotation protocol, high inter- or intra-rater variability, or inconsistencies in labeling the same entity across frames. Annotation variability may also arise from annotator fatigue, differences in clinical expertise, or a lack of annotation training and auditing procedures. Additionally, annotation quality may depend on task complexity, which may require multi-rater annotation and inter-/intra-rater agreement analysis. Even if annotations are consistent, they may still be unreliable, for example, if they are based on weak reference sources such as unreliable sensors or flawed clinical definitions. These inconsistencies and sources of unreliability can lead to models learning ambiguous or conflicting patterns, reduce the reliability of reported performance metrics, and make it difficult to interpret model failures or compare results across datasets or studies. | [nwoye2023cholectrack20, pangal2021guide, twinanda2016endonet, sharan2023mvhota, cartucho2024surgt, bano2024placental, mao2024pitsurgrt, radsch2023labelling, luca2022impact, demir2023deep, zisimopoulos2018deepphase, ghamsarian2023cataract] | Fig. 1a |
@@
| P1.6 | Poor data splitting | Data are not adequately split into training, validation, and test sets, leading to issues such as overfitting or data leakage. This could, for example, occur due to improper use of test splits for validation or model selection, non-stratified sampling, use of test splits for ablation studies, or the complete lack of a test set. A particularly harmful form of leakage can occur when data from the same patient is included in both training and test sets. This allows the model to leverage patient-specific characteristics, leading to overly optimistic performance estimates and reduced generalizability to new patients. The problem of data leakage is currently becoming even more severe with the emergence of generalist models trained on unknown cohorts, where the lack of transparency in the data composition makes proper separation and assessment especially challenging. | [funke2023metrics, liu2023skit, gao2021trans, wei2022boxpolyp, ahmad2024wavemamba, cervantes2021automatic, he2022empirical, nwoye2022data, kostiuchik2024surgical, jiang2021risk, abdulbaki2018surgical] | Extended Data Fig. 4 |
@@
| P2.1 | Mismatch of metrics and clinical needs | Metrics are used that do not reflect the underlying needs of the clinical use case or fail to capture the impact of the model’s performance on clinical decision-making. This often reflects a focus on technical performance rather than clinical relevance. For example, a high frame-wise accuracy in surgical phase recognition tasks may not be clinically meaningful if the model frequently misidentifies short but clinically relevant phases or critical transitions between surgical phases. Another example includes specific operational requirements such as real-time capabilities, which are often not considered in performance validation, despite being critical for clinical usability. Moreover, clinical needs are shaped not only by the task itself, but also by the intended role of the AI system, for example, whether it supports decision-making or acts autonomously. | [kankanamge2025artificial, hashimoto2024foundation, reinke2024understanding, kolbinger2024strategies, brand2022frame, bar2020impact] | Extended Data Fig. 5 |
@@
| P2.2 | Lack of metrics that assess temporal aspects | Metrics are used that do not assess algorithm properties specific to temporal data. For example, in surgical instrument segmentation, validating performance solely based on frame-wise Dice Similarity Score (DSC) may result in overlooking temporal inconsistencies, such as abrupt appearance or disappearance of instruments between consecutive frames. In addition to temporal coherence of predictions, metrics for evaluating real-time system behavior, such as latency or throughput, are often missing, despite their relevance in time-critical clinical settings. | [dergachyova2016automatic, boers2020improving, zhang2022retrieval, berlet2022surgical, sharan2023mvhota, luiten2021hota, kopf2021robust, funke2023metrics, brand2022frame] | Fig. 1b |
@@
| P2.3 | Inappropriate metric selection for handling annotation uncertainties | Annotation uncertainty is not adequately addressed in metric selection, which can result in misleading performance validation. This uncertainty may arise from inter- and intra-rater variability in subjective tasks, low visibility in surgical videos, or inconsistent labeling of ambiguous regions (e.g., tissue boundaries or transition moments) over time. For example, ambiguity in defining event boundaries - such as the exact moment a surgical phase transition occurs - may introduce inconsistencies in annotations due to inter- or intra-rater variability and missing exact ground truth. | [dergachyova2016automatic, funke2023metrics, berniker2021probabilistic, berlet2022surgical, nwoye2022data] | Extended Data Fig. 6 |
@@
| P2.4 | Lack of metric robustness across varying real-world conditions | The selected metrics lack robustness with respect to various real-world conditions such as data quality, acquisition settings, or clinical variability. For example, performance validation may be overly sensitive to changes in frame rate, image resolution, zooming, differences in surgical technique or intraoperative conditions, or annotation granularity. This includes common intraoperative phenomena such as smoke, motion blur, objects temporarily leaving the field of view, or changes in object size due to camera movement or zoom. These factors can distort metric behavior even when model predictions remain stable, and may substantially affect metric stability and reproducibility. | [nwoye2025surgitrack, czempiel2021opera, knoche2021susceptibility] | Extended Data Fig. 7 |
@@
| P2.5 | Non-standardized configuration and definition of metrics | Metrics without standardized configuration within a specific use case or definition are used, often without providing details on the concrete formula. For example, hyperparameters such as thresholds for surgical instrument detection, temporal tolerance ranges for transitions in phase recognition (e.g., how many frames of deviation are accepted as correct), or weighting of different error types may not be clearly defined or reported. In addition, even when the same metric is used, differences in spatial or temporal application, such as calculating accuracy only during annotated segments versus the full video, or validating phase recognition at different frame rates, can lead to non-comparable results and misinterpretation. | [funke2023metrics, gao2021trans, jin2021temporal, jin2017sv] | Extended Data Fig. 8a |

### v2

@@ 4.3. Experimental Design
Weighted phase-wise aggregation: Performance was first aggregated for each phase, and the resulting phase-level scores were then combined into a global score using clinician-defined weights reflecting the clinical relevance and complexity of each phase. For simplicity, we adopted a straightforward weight assignment strategy, in which phases deemed less relevant or complex were assigned a weight of 1 (phases 0, 4, 5, 6, 9, 12), intermediate phases were assigned 2 (phases 2 and 10), and highly relevant or complex phases received a weight of 3 (phases 1 and 8) [Maier-Hein et al., 2021].
Across all aggregation strategies, the aggregation operator can vary, e.g., it could be the mean, median, or other percentiles. In line with the original challenge, we used the 5th percentile to reflect worst-case performance. To assess the agreement between the default (frame-wise) ranking scheme and alternative aggregation strategies, we calculated Kendall’s tau [Kendall, 1938], a rank correlation coefficient, quantifying the similarity between two orderings (Kendall’s tau = 1: perfect agreement, Kendall’s tau = 0: no association, Kendall’s tau = -1: complete disagreement).
@@ 5. Extended Data
6. Data availability
The experiments in this study were conducted using publicly available datasets, namely the RobustMIS 2019 challenge data [Roß et al., 2021] and the CholecT45 data [Nwoye et al., 2023a]. Individual challenge results (i.e., metric scores per algorithm) are not publicly available, as participant-level consent for data sharing was not obtained.
7. Code availability
Source code for the three experiments is available at https://github.com/IMSY-DKFZ/metrics-reloaded-surgery-pitfalls.
@@ 8. Acknowledgements
A.R.: This work was initiated by the Helmholtz Association of German Research Centers in the scope of the Helmholtz Imaging Incubator (HI). L.M.-H.: This project was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (NEURAL SPICING, 101002198), the National Center for Tumor Diseases (NCT), Heidelberg’s Surgical Oncology Program, the German Cancer Research Center (DKFZ). E.C.: This publication was further supported through state funds approved by the State Parliament of Baden-Württemberg for the Innovation Campus Health + Life Science Alliance Heidelberg Mannheim. O.C.: The research leading to these results has received funding from the French government under management of Agence Nationale de la Recherche as part of the “France 2030” program (reference ANR-23-IACL-0008, project PRAIRIE-PSAI), as part of the ”Investissements d’avenir” program (reference ANR-19-P3IA-0001, project PRAIRIE 3IA Institute and reference ANR-10-IAIHU-06, project Agence Nationale de la Recherche-10-IA Institut Hospitalo-Universitaire-6) and from the European Union’s Horizon Europe Framework Programme (grant number 101136607, project CLARA). R.D.: This project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101092646. D.D. is funded by NIH K23 EB034110. Q.D.: This work described in this paper was partially supported by a grant from the ANR/RGC Joint Research Scheme sponsored by the Research Grants Council of the Hong Kong Special Administrative Region, China and the French National Research Agency (Project No. A-CUHK402/23). G.F. is supported by Canada Research Chair in Computer-Assisted Surgery. S.Gia. is supported by the Royal Society URF R 201014. T.H. is a Consolidator Researcher, receiving financial support from the Distinguished Researcher program of Obuda University. His work has been partially supported by ACMIT (Austrian Center for Medical Innovation and Technology), which is funded within the scope of the COMET (Competence Centers for Excellent Technologies) program of the Austrian Government. T.H. is partially supported by Project 2024-1.2.3-HU-RIZONT-00069, implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the 2024-1.2.3-HU-RIZONT funding scheme. D.A.H. is supported by a grant from the American Surgical Association Foundation. F.R.K. receives support from the German Cancer Research Center (CoBot 2.0), the Joachim Herz Foundation (Add-On Fellowship for Interdisciplinary Life Science), the Central Indiana Corporate Partnership AnalytiXIN Initiative, the Evan and Sue Ann Werling Pancreatic Cancer Research Fund, and the Indiana Clinical and Translational Sciences Institute (EPAR4157) funded, in part, by Grant Number UM1TR004402 from the National Institutes of Health, National Center for Advancing Translational Sciences, Clinical and Translational Sciences Award. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. J.L.L. received funding by the Swiss National Science Foundation (P5R5PM 21766), Novartis Foundation for medical-biological Research (#23C162), and the Vontobel Foundation (0867/2024). H.J.M. is supported by the NIHR UCLH/UCL Biomedical Research Centre. N.P.: This work has received funding from the European Union (ERC, CompSURG, 101088553, PI: Nicolas Padoy). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work has also been supported by French state funds managed within the Plan Investissements d’Avenir by the ANR under reference ANR- 10-IAHU-02 (IHU Strasbourg). H.R.: Hong Kong Research Grants Council Collaborative Research Fund under Grant CRF-C4026-21G. S.S. is supported by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence ”Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Dresden University of Technology. D.S. is supported by the Royal Academy of Engineering Chair in Emerging Technologies. M.W.: This work has been funded by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy—EXC 2050/1—Project ID 390696704—Cluster of Excellence ”Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of TUD Dresden University of Technology and by the German Federal Ministry of Research, Technology, and Space within the “Surgical AI Hub Germany” project (grant number BMBF 02K223A110).
@@ 9. Conflicts of Interest
M.S.A. is a speaker of Medtronic and advisor of Johnson & Johnson. D.A. is a shareholder and has a leadership role in Scialytics. N.W.C.: Nil. O.C. reports having received consulting fees from Therapanacea (2022-2024). OC reports that other principal investigators affiliated to the team which he co-leads have received grants (paid to the institution) from Sanofi (2020-2022) and Biogen (2022-2023). OC reports that his spouse was an employee of myBrainTechnologies (2015-2023) and is an employee of DiamPark. G.F. declares the following conflict of interest: Johns Hopkins University, Harvard Brigham and Women’s Hospital. S.G. declares the following competing financial interests: he has or has had consulting relationships with Una Health GmbH, Lindus Health Ltd., Flo Ltd, ICURA ApS, Rock Health Inc., Thymia Ltd., FORUM Institut für Management GmbH, High-Tech Gründerfonds Management GmbH, DG SANTE, Prova Health Ltd, haleon plc and Ada Health GmbH and holds share options in Ada Health GmbH. T.P.G. is the Founder of SST. D.A.H. is a consultant for Medtronic. A.J. is an employee of Intuitive. F.R.K. declares advisory roles for Radical Healthcare, USA; and the Surgical Data Science Collective, USA, and has received research funding from Novartis. H.J.M. is employed by and holds shares in Panda Surgical Limited. P.M. is a co-founder and shareholder of Scialytics. H.N. is affiliated with Proximie Ltd, London, UK. N.P. is co-founder and own shares in Scialytics SAS. N.R. is an employee at NVIDIA. M.M.R. is now working at SAP Fioneer. D.S.: Medtronic, Odin Vision, Panda Surgical, EnAcuity, Helico Medical, Uncovr, Vope Medical. M.W. has received honoraria from KARL STORZ SE & Co KG.
@@
| SUPPL. NOTE 3 | Practical best practices | SUPPL. NOTE 3 |
| SUPPL. NOTE 4 | Results of the systematic review per pitfall | SUPPL. NOTE 4 |
| SUPPL. NOTE 5 | Comparison of validation practices across MICCAI 2023 and 2025 | SUPPL. NOTE 5 |
@@
| P1.1 | Non-representativeness, low quality, and low relevance of data | The data used for training and/or testing does not accurately reflect the intended real-world use case or lacks clinical relevance. This may, for example, result from geographical imbalance, a lack of diversity in patient demographics, regional differences in surgical standards, or limited or outdated variations in surgical techniques. Additionally, data selection criteria may exclude challenging but clinically relevant cases, such as poor camera quality, low lighting conditions, challenging procedures, or imaging artifacts, which are crucial for assessing model robustness in realistic scenarios. The inclusion of irrelevant data, such as out-of-body frames or unintentionally recorded segments, may also dilute the training signal or introduce misleading patterns. The use of simulated or experimental data may further reduce real-world applicability, while proxy data can misalign validation with clinical objectives. Variability in video quality and preprocessing protocols across institutions may further compromise data fidelity and model reproducibility. In addition, unobserved operator-related heterogeneity, such as multiple surgeons with different experience levels (e.g., trainees and attending surgeons) contributing to the same procedure, particularly in training centers, may introduce additional variance if operator roles are not annotated or reported. | [Godau et al., 2023, Kolbinger et al., 2024b, Koch et al., 2024, Sellner et al., 2023, Kilim et al., 2022, Castro et al., 2020, Subasri et al., 2025, Lavanchy et al., 2024, Kassem et al., 2022, Mascagni et al., 2022, Kong et al., 2021, Junger et al., 2022, Bar et al., 2020, Paranjape et al., 2023, Engelhardt et al., 2018, Sahu et al., 2021, Huaulmé et al., 2023, Kolbinger et al., 2023, 2024a, Ding et al., 2022, Childers and Maggard-Gibbons, 2021, Jogan et al., 2024] | Extended Data Fig. 1 |
@@
| P1.2 | Limited sample size/test cohorts, or data imbalance | The test data may be limited due to small sample size and/or the absence of validation on independent datasets, including those from different institutions, time periods, or populations (external validation), or from real-world clinical use after deployment (post-deployment validation). Limited data can mislead model development (e.g., algorithm selection), increase the risk of overfitting, or yield unreliable model performance validation, making it difficult to draw meaningful conclusions about the model’s real-world performance. These issues are amplified in the presence of class imbalance, where rare classes may appear in only one subset or fold, or be entirely absent, leading to inconsistent model performance depending on the specific validation or test set used. | [Bar et al., 2020, Kostiuchik et al., 2024, Yamlahi et al., 2023, Demir et al., 2023, Killeen et al., 2023, Maqbool et al., 2020, Cheng et al., 2023] | Extended Data Fig. 2 |
@@
| P1.3 | Existence of spurious correlations within data | Spurious correlations are statistical patterns in the data that do not reflect robust or generalizable associations with the target task. They can arise from biases in data collection, labeling, or contextual factors unrelated to the underlying clinical objective, and often reflect dataset-specific artifacts that fail to generalize. This may lead to short-cut learning, where a model exploits irrelevant patterns, for example, achieving seemingly good performance by associating certain surgical outcomes with specific recording conditions, such as particular camera types used at different hospitals, or inferring the presence of instruments from irrelevant features such as glove color, if such features are biased within the dataset. | [Yuan et al., 2024, Banerjee, 2024, Mahmood et al., 2021, Geirhos et al., 2020, Dammu and Shah, 2023, Bykov et al., 2023] | Extended Data Fig. 3a |
@@
| P1.4 | Incomplete annotation or missing contextual information | Annotations of the data are – intentionally or unintentionally – incomplete, leading to gaps in the information available for reliable validation. This may include missing metadata, partially annotated video sequences, the omission of specific objects, events, or surgical phases, or insufficient detail for the intended task. Missing metadata, such as age, for example, may lead to undetected confounders and thus biased models. Partially annotated video data may render the application of temporal metrics (e.g., for assessing continuity or stability across frames) infeasible and thus limit the ability to assess performance over time, such as continuity or stability across video frames. While some approaches may use partial annotations intentionally (e.g., in weak supervision), unmanaged or undocumented incompleteness may still compromise reliability and validation validity. The omission of relevant entities can result in inaccurate performance estimates, for example, underestimating sensitivity when relevant instances are not labeled, and may give a misleading impression of model accuracy and clinical utility. | [Pangal et al., 2021, Ward et al., 2021, Tajbakhsh et al., 2020, Twinanda et al., 2016, Funke et al., 2019, Peng et al., 2024, Nyangoh Timoh et al., 2023, Murali et al., 2023] | Extended Data Fig. 3b |
@@
| P1.5 | Unreliable or inconsistent annotation | Unreliable or inconsistent annotations can undermine the validity of model validation by introducing ambiguity, bias, or uncontrolled variability. This may result from the lack or insufficiency of a standardized annotation protocol, high inter- or intra-rater variability, or inconsistencies in labeling the same entity across frames. Annotation variability may also arise from annotator fatigue, differences in clinical expertise, or a lack of annotation training and auditing procedures. Additionally, annotation quality may depend on task complexity, which may require multi-rater annotation and inter-/intra-rater agreement analysis. Even if annotations are consistent, they may still be unreliable, for example, if they are based on weak reference sources such as unreliable sensors or flawed clinical definitions. These inconsistencies and sources of unreliability can lead to models learning ambiguous or conflicting patterns, reduce the reliability of reported performance metrics, and make it difficult to interpret model failures or compare results across datasets or studies. | [Nwoye et al., 2023b, Pangal et al., 2021, Twinanda et al., 2016, Sharan et al., 2023, Cartucho et al., 2024, Bano et al., 2024, Mao et al., 2024, Rädsch et al., 2023, Luca et al., 2022, Demir et al., 2023, Zisimopoulos et al., 2018, Ghamsarian et al., 2023] | Fig. 1a |
@@
| P1.6 | Poor data splitting | Data are not adequately split into training, validation, and test sets, leading to issues such as overfitting or data leakage. This could, for example, occur due to improper use of test splits for validation or model selection, non-stratified sampling, use of test splits for ablation studies, or the complete lack of a test set. A particularly harmful form of leakage can occur when data from the same patient is included in both training and test sets. This allows the model to leverage patient-specific characteristics, leading to overly optimistic performance estimates and reduced generalizability to new patients. The problem of data leakage is currently becoming even more severe with the emergence of generalist models trained on unknown cohorts, where the lack of transparency in the data composition makes proper separation and assessment especially challenging. | [Funke et al., 2023, Liu et al., 2023, Gao et al., 2021, Wei et al., 2022, Ahmad et al., 2024, Cervantes-Sanchez et al., 2021, He et al., 2022, Nwoye and Padoy, 2022, Kostiuchik et al., 2024, Jiang et al., 2021, Abdulbaki Alshirbaji et al., 2018] | Extended Data Fig. 4 |
@@
| P2.1 | Mismatch of metrics and clinical needs | The used metrics do not reflect the underlying needs of the clinical use case or fail to capture the impact of the model’s performance on clinical decision-making. This often reflects a focus on technical performance rather than clinical relevance. For example, a high frame-wise accuracy in surgical phase recognition tasks may not be clinically meaningful in case the model frequently misidentifies short but clinically relevant phases or critical transitions between surgical phases. Another example includes specific operational requirements such as real-time capabilities, which are often not considered in performance validation, despite being critical for clinical usability. Moreover, clinical needs are shaped not only by the task itself, but also by the intended role of the AI system, for example, whether it supports decision-making or acts autonomously. | [Kankanamge et al., 2025, Hashimoto et al., 2024, Reinke et al., 2024, Kolbinger et al., 2024b, Brand et al., 2022, Bar et al., 2020] | Extended Data Fig. 5 |
@@
| P2.2 | Lack of metrics that assess temporal aspects | The used metrics do not assess algorithm properties specific to temporal data. For example, in surgical instrument segmentation, validating performance solely based on frame-wise Dice Similarity Score (DSC) may result in overlooking temporal inconsistencies, such as abrupt appearance or disappearance of instruments between consecutive frames. In addition to temporal coherence of predictions, metrics for evaluating real-time system behavior, such as latency or throughput, are often missing, despite their relevance in time-critical clinical settings. | [Dergachyova et al., 2016, Boers et al., 2020, Zhang et al., 2022, Berlet et al., 2022, Sharan et al., 2023, Luiten et al., 2021, Kopf et al., 2021, Funke et al., 2023, Brand et al., 2022] | Fig. 1b |
@@
| P2.3 | Inappropriate metric selection for handling annotation uncertainties | Annotation uncertainty is not adequately addressed in metric selection, which can result in misleading performance validation. This uncertainty may arise from inter- and intra-rater variability in subjective tasks, low visibility in surgical videos, or inconsistent labeling of ambiguous regions (e.g., tissue boundaries or transition moments) over time. For example, ambiguity in defining event boundaries - such as the exact moment a surgical phase transition occurs - may introduce inconsistencies in annotations due to inter- or intra-rater variability and missing exact ground truth. | [Dergachyova et al., 2016, Funke et al., 2023, Berniker et al., 2021, Berlet et al., 2022, Nwoye and Padoy, 2022] | Extended Data Fig. 6 |
@@
| P2.4 | Lack of metric robustness across varying real-world conditions | The selected metrics lack robustness with respect to various real-world conditions such as data quality, acquisition settings, or clinical variability. For example, performance validation may be overly sensitive to changes in frame rate, image resolution, zooming, differences in surgical technique or intraoperative conditions, or annotation granularity. This includes common intraoperative phenomena such as smoke, motion blur, objects temporarily leaving the field of view, or changes in object size due to camera movement or zoom. These factors can distort metric behavior even when model predictions remain stable, and may substantially affect metric stability and reproducibility. | [Nwoye and Padoy, 2025, Czempiel et al., 2021, Knoche et al., 2021] | Extended Data Fig. 7 |
@@
| P2.5 | Non-standardized configuration and definition of metrics | Metrics without standardized configuration within a specific use case or definition are used, often without providing details on the concrete formula. For example, hyperparameters such as thresholds for surgical instrument detection, temporal tolerance ranges for transitions in phase recognition (e.g., how many frames of deviation are accepted as correct), or weighting of different error types may not be clearly defined or reported. In addition, even when the same metric is used, differences in spatial or temporal applications, such as calculating accuracy only during annotated segments versus the full video, or validating phase recognition at different frame rates, can lead to non-comparable results and misinterpretation. | [Funke et al., 2023, Gao et al., 2021, Jin et al., 2021, 2017] | Extended Data Fig. 8a |